Consider the function f(x) = 5 ; for x < 0 0, for x = 0 5, for x > 0

Show that f has no limit as x approaches 0


1
Expert's answer
2021-08-12T13:20:18-0400
"f(x) = \begin{cases}
 5 &\text{for } x<0 \\
 0 &\text{for } x=0 \\
 5 &\text{for } x>0 \\
\end{cases}"

"\lim\limits_{x\to 0^-}f(x)=5"

"\lim\limits_{x\to 0^+}f(x)=5"

Since "\lim\limits_{x\to 0^-}f(x)=5=\lim\limits_{x\to 0^+}f(x)," then "\lim\limits_{x\to 0}f(x)" exists, and


"\lim\limits_{x\to 0}f(x)=5"